Shift Summary for Owl shift, Nov 5, 1999

Yee Bob Hsiung (hsiung@fnal.gov)
Fri, 5 Nov 1999 07:25:39 -0600

Shift Summary for Owl shift, Nov 5, 1999

On Shift: Sasha Glazov, Yee Bob Hsiung

Hours of beam delivered at all intensities: 2:30
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 1:40
Hours of special runs (laser scans, muon runs...):0:25

Accounting of lost time: TeVatron startup problems, quenches, PS ramp, run
changes.

=========================================================================

Runs ended this shift:
Run 14906 - Beam_Muon_4E12, 13M events.
Run 14907 - KTeV_Ped. for Update peds.
Run 14909 - Beam_799, still in progress. Debugging run

Tapes brought to FCC this shift: none

=========================================================================

Overview of shift:

Loooooong wait for Tevatron startup, then first beam arrived at 4:23am
spill 379, run 14906. Intensity <=4E12 for our startup and debugging.
First E799 hadron run 14909 has no accidentals, no BA1-L(+) trig source,
problem in ytf and kq or HCC hot bit (affecting L2 yield, especially for
4-track and Dalitz)?!?!?! Minor problem for RC8A trig source too.
Stay tune for next shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14906 1:05 Tuning for Tevatron in progress, beam in Tevatron,
1.5E12 to SY dump. Still waiting for beam to us.
several quenches in Tevatron due P1 transfer line.
379 4:23 1st muon beam! Wooooooaaaaaaa! ~3.5E12
page BobT.
4:40 stop muon run.
change to normal E799 hadron beam config.
ramp up NM4AN and open beam stop.
14907 4:47 Pedestal run.
4:53 stop run and update pedestal and DC alignment
constants.
14909 5:18 finally run started after problem init kq.
checking various detector alives or not:
ACC 90-degree Monitor not counting (though
ACC1-3 are all alive). BA1-L not counting,
but ADC looks alright.
L2 rate does not look right, problems with
ytf and kq?! investigating and paged Jason.
HCC hot bit (from Etot)?!
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Continue debugging the problems mentioned above.